Sin(A) = -3/5 with A in quadrant III. Find tan (2A).
Answer by lwsshak3(6505)

Sin(A) = -3/5 with A in quadrant III. Find tan (2A).
**
You are working with a right triangle in quadrant III where the opposite side=3, and hypotenuse=5. By the Pythagorean Theorem, the adjacent side=4. Tan(A)=-3/4
Identity:Tan(2A)=2tan(A)/(1-tan^2(A))
Tan(2A)=2(-3/4)/(1-(-3/4)^2)
Tan(2A)=-(3/2)/(7/16)=48/14=-24/7